What amount must Harold save each month for the next 4 years if he wants to accumulate $40,000 and he can earn an annual rate of 9 percent?  His first deposit will be one month from today. (Choose the closest answer)
Answer #1

pmt x ((1+r)^n-1))/r Future value = PMT = FV+ ((1+r)^n-1))/r Save each month PMT =|$695.40 [40,000/(((1.0075^48)-1)/0.0075)]